OpenStudy (anonymous):

please help What distance is traveled by a satellite 1200 km from the center of the Earth in 40pi € rotations?

OpenStudy (anonymous):

these are the choices A) 48,000 B) 150,796 C) 225,000 D) 250,435

OpenStudy (anonymous):

@hartnn turns out that the " l " symbol is actually Pi (3.1415)

hartnn (hartnn):

in one rotation, satellite will cover one circumference distance. right ?

OpenStudy (anonymous):

so 1200 is the radius r

OpenStudy (anonymous):

circum. =2 Pi r so 2 x 3.1415 x 1200

hartnn (hartnn):

in 40pi rotations, distance covered = 40pi times circumference \(D =40\pi (2\pi r) \)

hartnn (hartnn):

none of the option match...

OpenStudy (anonymous):

yes that is the problem :(

hartnn (hartnn):

you sure about being "pi" ?

OpenStudy (anonymous):

yes the teacher replied back and said it is pi

OpenStudy (anonymous):

is there any other way to solve this ?

hartnn (hartnn):

\(40\pi r = 40 \pi (1200) = 150796\)

OpenStudy (anonymous):

thats B , in the choices

OpenStudy (anonymous):

i think this is the answer , right ?

hartnn (hartnn):

its like saying, that in 2pi rotations, distance covered is 'r' so in 40pi rotations, distance covered will be 40pi r does not seem quite correct...logically.

hartnn (hartnn):

Logic says, distance covered in one rotation = 2pir

OpenStudy (anonymous):

i see but it is the closest in the choices

OpenStudy (anonymous):

yes in one rotation so its covering the circum. of the circle

hartnn (hartnn):

given those choices, preferable to probably pick up B) only . if and when you get an explanation for this, share with us. @ganeshie8 might get some sense out of it...

hartnn (hartnn):

probably \(\Large 2\pi\) rotations would mean \(\Large 1 \) round around the earth. so , 40pi rotations would mean only 40pi/2pi = \(\Large 20\) rounds of earth giving 20* (2pi r) = 150796 km distance

OpenStudy (anonymous):

yes I wanted to post in the yesterday's post but unfortunately its closed. so as soon as I'll get the explanation . I will post a new post and share it with both of you.

OpenStudy (anonymous):

probably , thats the best and closest explanation so far , thank you so much for your effort @hartnn

OpenStudy (anonymous):

I'll share the explanation as soon as I'll get it

hartnn (hartnn):

thanks :) and welcome ^_^